Solve the differential equation:



a) yk - y(k-1) + 2y(k-2) = k² + 5k




b) y(k+2) - 4y(k+1) + yk = 3k +2^k




In LHS yk , y(k-1) and so on is not in multiply but the k part is written in down to y( like yk is not y×k but the k is written in right down of y).

1
Expert's answer
2021-12-15T16:21:47-0500

a)

"y_k - y_{k-1} + 2y_{k-2} = k\u00b2 + 5k"


charasteristic equation:

"r^2-r+2=0"


"r=\frac{1\pm i \sqrt 7}{2}"


"y_h=r^k(acos(k\theta)+bsin(k\theta))"


"r=\sqrt{(1/2)^2+(\sqrt 7/2)^2}=\sqrt 2"

"\theta=arccos(1/(2\sqrt 2))=1.2" rad


"y_h=(\sqrt 2)^k(acos(1.2k)+bsin(1.2k))"


"y_t=Ak^2+Bk+C"


"Ak^2+Bk+C+A(k-1)^2+B(k-1)+C+2A(k-2)^2+2B(k-2)+2C="

"=k^2+5k"


"4A=1\implies A=1/4"

"4B-10A=5\implies B=1.875"

"4C+9A-5B=0\implies C=1.78"


"y_k=(\sqrt 2)^k(acos(1.2k)+bsin(1.2k))+0.25k^2+1.875k+1.78"


b)

"y_{k+2} - 4y_{k+1} + y_k = 3k +2^k"


charasteristic equation:

"r^2-4r+1=0"


"r=2\pm \sqrt 3"

"y_h=a(2-\sqrt 3)^k+b(2+\sqrt 3)^k"


"y_{t1}=Ak+B"

"A(k+2)+B-4(A(k+1)+B)+Ak+B=3k"

"-2A=3\implies A=-1.5"

"-2A-2B=0\implies B=1.5"


"y_{t2}=A2^k"

"A2^{k+2}-4A2^{k+1}+A2^k=2^k"

"4A-8A+A=1"

"A=-1\/3"


"y_k=a(2-\sqrt 3)^k+b(2+\sqrt 3)^k-1.5k+1.5-2^k/3"